The profit on a watch is given by P=x^2-13x-80 and where x is the number of watches sold her day. How many watches were sold on a day when there was a $50 loss?
Answer by nerdybill(7384)

P=x^2-13x-80
How many watches were sold on a day when there was a $50 loss?
.
This means P is -50.
Plug it in and solve for x:
P=x^2-13x-80
-50=x^2-13x-80
0=x^2-13x-30
factoring:
0=(x-15)(x+2)
.
x = {-2, 15}
.
We can toss out the negative answer so we get:
x = 15 watches